OpenStudy (anonymous):

One-third of eighth graders in Wilson Middle School own a cell phone. If 75 students own cell phones, how many eighth graders are in Wilson Middle School? Choose the correct equation and solution for this situation.

OpenStudy (anonymous):

1/3 + x = 75; x = 225 3x = 75; x = 225 1/3x = 75; x = 225 75/x = 1/3; x= 225

OpenStudy (anonymous):

well, 1/3 or eighth graders own cell phones, and that's 75 students, wouldn't you have had to take 1/3 of the number of total students to get that 75?

OpenStudy (anonymous):

so, what does the x stand for in the prob?

OpenStudy (anonymous):

so, what does the x stand for in the prob?